n+5=-1, the first part has a denominator of -16
Found 2 solutions by solver91311, rfer:
Answer by solver91311(24713)



Multiply both sides by -16. Add -5 to both sides.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by rfer(16322)

I hope I understand
(n+5)/-16=-1
n+5=-1*-16
n+5=+16
n=16-5
n=11